### v1

# Nontrivial twisted bilayer chiral excitonic systems: Förster coupling and related Hall effect

Abstract
In bilayers composed of monolayers of two-dimensional semiconductors with arbitrary twisting angles, an effective excitonic system can be established based on the valley excitons of each layer. Instead of direct interlayer hopping, a pronounced Förster coupling connects intralayer excitons in different layers, which modifies the energy of the monolayer excitonic system. Surprisingly, two new types of Hall effect recently discovered in electric systems-the time-reversal even charge Hall effect (TREHE) and the crossed nonlinear dynamical Hall effect (CNDHE)-can occur in these systems due to the chiral structure induced by the twist of two layers. We illustrate this point by employing a general theoretical framework to establish the effective excitonic Hamiltonians in twisted homobilayer systems composed of transition metal dichalcogenides (TMD) and black phosphorus (BP), calculating the Hall effect in these systems as examples. Regarding the TREHE, our calculations show that it depends on the anisotropy of excitons, resulting in a negligible Hall conductivity for TMD without strain due to the isotropic excitons, but a large Hall conductivity for TMD with chiral symmetric excitons induced by strain. Meanwhile, BP exhibits a finite Hall conductivity due to its high anisotropy. As for the CNDHE, only the specific value of twisted angles can induce a significant Hall conductivity in BP and TMD with strain. This work uncovers nontrivial physical properties of twisted bilayer chiral excitonic systems and provides new opportunities for valley exciton optoelectronics in the layer degree of freedom.
The study of emerging layered two-dimensional (2D) semiconductors has increasingly become one of the most crucial fields in condensed matter physics [1, 2, 3, 4, 5, 6]. Several of these materials can serve as a reliable platform to explore exciton physics due to strong light-matter interactions [2, 3, 4]. In transition metal dichalcogenides (TMDs), tightly bound Wannier excitons can emerge around the degenerate ±$\pm K$ valleys at the corners of the Brillouin zone (BZ), with optical selection rules for manipulating the valley pseudospin [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The small Bohr radius $\sim O(1)$ nm [17, 18, 19] indicates a significant electron-hole (e-h) Coulomb exchange at finite center-of-mass (COM) momentums [13, 20, 21, 22]. This coupling between the exciton’s valley pseudospin and COM degrees of freedom results in the splitting of the exciton dispersion into two branches having linearly polarized optical dipoles, longitudinal ($L$) and transverse ($T$) to exciton momentum, respectively [13, 21, 23, 24, 25, 26]. Recently, researchers have studied the exotic properties of ground-state excitons in monolayer TMDs on patterned substrates [27, 28] due to the massless $L$ branch’s sensitive dependence on the surrounding dielectric [29, 30]. Also, the $C_{3}$ symmetry of monolayer TMDs can be broken by applying the strain into it, which leads to a non-zero contribution of short-range e-h exchange between excitons in different valleys, causing linearly dispersing Dirac saddle points in the light cone [13, 21].
Otherwise, in highly anisotropic materials such as the black phosphorus (BP), which is a potentially ideal platform for high-performance devices like a field-effect transistor [31] since monolayer BP is a semiconductor with a predicted direct bandgap of $1\sim 2\,\mathrm{eV}$ at the $\Gamma$ point in the BZ [32, 33, 34]. The highly anisotropic band structure of the lowest conduction band and highest valence band [32, 33, 34] leads to anisotropic valley excitons [33, 35] corresponding with almost linearly polarized optical dipole instead of circularly polarized one for ±$\pm K$ isotropic valley excitons in monolayer TMDs [21, 22, 23].
In contrast to the extensively studied electrical behaviors in twisted bilayer and multi-layer systems [2, 5, 6, 7], or excitonic systems that rely on the interaction between interlayer and intralayer excitons [3, 4, 21, 37], the effective system related to Förster coupling [38], specifically the Coulomb exchange between electron-hole (e-h) pairs, which can transfer an exciton or e-h pair non-locally between intralayer excitons in different layers, has seldom been considered. This interaction enables out-of-plane propagation of excitons in a stack of layers, and its long-range nature allows it to work even when charge hopping is completely quenched by spacer layers or crystalline misalignment. Also, previous studies have proved that cross-dimensional valley excitons can arise from this coupling in arbitrarily twisted stacks of monolayer semiconductors [39].
Just as the anomalous Hall effect for electrons that results from the Berry curvature in momentum space [40], the exciton Hall effect can also arise due to the exciton Berry curvature [41, 42]. Recently, two new types of Hall effect have been discovered in twisted bilayer chiral electric systems, i.e., the time-reversal even charge Hall effect (TREHE) based on the spectral chiral symmetry [43] and the crossed nonlinear dynamical Hall effect (CNDHE) in twisted bilayers, induced by out-of-plane alternating-current (AC) electric fields [44, 45]. These findings uncover novel Hall physics in chiral structures and pave the way for layertronics, a promising avenue for exploring the quantum nature of the layer degree of freedom and discovering exciting effects. The question that naturally arises is whether these effects can also occur in twisted bilayer chiral excitonic systems and be experimentally observed.
In this paper, we investigate the properties of twisted bilayer chiral excitonic systems that arise from intralayer excitons in each layer and the Förster coupling connecting them across different layers. Using twisted homobilayer TMDs and BP as examples, we build a general method to derive effective excitonic Hamiltonians from electron-hole Coulomb exchange interactions and demonstrate that two newly discovered Hall effects, i.e., the TREHE and the CNDHE, in electronic systems also exist in excitonic systems. Our calculations reveal that the twist-induced chiral structure, material anisotropy, and twist angle significantly influence these effects. These findings reveal nontrivial properties and present new opportunities for valley exciton optoelectronics.
Figure 1: (color online) (a) Left panel: Schematic of an arbitrarily twisted homobilayer BPs, where hBN spacers quench the charge hopping. Right panel: Illustration of electron-hole Coulomb exchange between valley excitons in the same or different layers, showing approximate band edges near high-symmetry points. (b) Similar plot for twisted homobilayer TMDs with strain (applied along $x$ direction with strain extent $\epsilon_{st}\sim 1\%$ in the bottom layer). (c) Left panel: Excitonic dispersions for the effective excitonic Hamiltonian (26) of BP with different twisted angles $\theta$. Right panel: Anisotropic polarization distribution of the optical dipole for two bands in momentum space. (d) Similar plots for the effective excitonic Hamiltonian (17) of TMD with and without strain in untwisted situation. For clarity, only the polarization distribution $p_{n}$ of the optical dipole for the marked band $\varepsilon_{n}$ are shown.
I Effective excitonic Hamiltonians of twisted homobilayer TMDs and BP
As previously stated, the effective excitonic Hamiltonian in twisted bilayer structures is primarily based on intralayer valley excitons in the two layers. In the case of monolayer situations under the low-energy approximation, there are typically at least two non-equivalent valleys connected by rotational or time-reversal symmetry in the first BZ, such as the ±$\pm K$ valley in monolayer TMDs [2, 7] (Fig. 1(b)), resulting in a $4\times 4$ effective Hamiltonian for intralayer excitons, as discussed below. However, when only a single valley is present, such as the $\Gamma$ point in monolayer BP (Fig. 1(a)), the Hamiltonian reduces to a simple $2\times 2$ matrix. In the following, we present a method for obtaining the effective excitonic Hamiltonian through a general calculation of the e-h Coulomb exchange interaction, and use it to acquire the excitonic Hamiltonian for these two materials.
I.1 General description the effective excitonic Hamiltonian
The construction of this Hamiltonian relies on two types of interaction: the e-h exchange interaction between intralayer excitons of the same or different valleys in the same layer and the Förster coupling that connects excitons in different layers, as illustrated in Fig. 1. For a specific layer $l$ ($l=t,b$ in this paper, $t/b$ means top/bottom layer, as shown in Fig. 1(a)), the basis of valley excitons can be expressed as $\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$, $\alpha,\beta$ means two different valley indexes with the in-plane COM momentum $\boldsymbol{k}=\left(k\cos\varphi,k\sin\varphi\right)$, and generally the effective excitonic Hamiltonian in this situation can be expressed as
$H_{\mathrm{ex}}=\left(\begin{array}[]{cc}H_{k}^{b}&0\\ 0&H_{k}^{t}\end{array}\right)+\left(\begin{array}[]{cc}H_{\mathrm{intra}}^{b}&H_{\mathrm{inter}}^{b,t}\\ H_{\mathrm{inter}}^{t,b}&H_{\mathrm{intra}}^{t}\end{array}\right).$ (1)
The first term is the kinetic energy of excitons in different layers, which has the form
$H_{k}^{l}=\left(\begin{array}[]{cc}\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex,\alpha}^{\gamma,l}\right)&0\\ 0&\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex,\beta}^{\gamma,l}\right)\end{array}\right),$ (2)
with $m_{ex,\alpha}^{\gamma,l},\gamma=x,y$ is the exciton mass in $l$ layer, depending on the twisted angle $\theta_{l}$ of each layer (see Supplementary [36] for more details).
$\left(\begin{array}[]{c}\left\langle l,\boldsymbol{k}\right|_{\alpha}\\ \left\langle l,\boldsymbol{k}\right|_{\beta}\end{array}\right)H_{\mathrm{intra}}^{l}\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$ (5)
$\equiv$ $\left(\begin{array}[]{cc}J_{\alpha,\alpha}^{l}&J_{\alpha,\beta}^{l}\\ J_{\beta,\alpha}^{l}&J_{\beta,\beta}^{l}\end{array}\right),$ (8)
represents the e-h intra- or intervalley exchange between the excitons in the same or different valleys in the layer $l$, where
$J_{\lambda,\lambda^{\prime}}^{l}$$\approx$ $\psi_{\lambda}^{l}\left(0\right)\left[\psi_{\lambda^{\prime}}^{l}\left(0\right)\right]^{\ast}\frac{V\left(\boldsymbol{k}\right)}{4}\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda}^{+}\right)\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda^{\prime}}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda^{\prime}}^{+}\right)^{\ast},$
with $k_{\pm}=k_{x}\pm ik_{y}$, $d_{cv,\lambda}^{\pm}=d_{cv,\lambda}^{x}\pm id_{cv,\lambda}^{y}$. $d_{cv,\lambda/\lambda^{\prime}}^{\gamma},\lambda,\lambda^{\prime}=\alpha,\beta$ is the optical transition dipole in $\gamma$ direction, between conduction ($c$) and valence ($v$) band edges of $\lambda$ or $\lambda^{\prime}$ valley. $\psi_{\lambda}^{l}\left(0\right)\sim 1/\sqrt{a_{B,\lambda}^{x,l}a_{B,\lambda}^{y,l}}$ with the exciton Bohr radius $a_{B,\lambda}^{\gamma,l}$ for $\lambda$ valley can be seen as the square root of the probability for electron and hole to overlap in an exciton. $V\left(\boldsymbol{k}\right)=2\pi e^{2}/\left(\epsilon k\right)$ is the unscreened form for the Coulomb potential with $\epsilon\equiv 4\pi\epsilon_{0}\epsilon_{r}$ [13, 21, 23].
$\left(\begin{array}[]{c}\left\langle l,\boldsymbol{k}\right|_{\alpha}\\ \left\langle l,\boldsymbol{k}\right|_{\beta}\end{array}\right)H_{\mathrm{inter}}^{l,l^{\prime}}\left(\left|l^{\prime},\boldsymbol{k}\right\rangle_{\alpha},\left|l^{\prime},\boldsymbol{k}\right\rangle_{\beta}\right)$ (12)
$\equiv$ $\left(\begin{array}[]{cc}J_{\alpha,\alpha}^{l,l^{\prime}}&J_{\alpha,\beta}^{l,l^{\prime}}\\ J_{\beta,\alpha}^{l,l^{\prime}}&J_{\beta,\beta}^{l,l^{\prime}}\end{array}\right),$ (15)
shows the Förster coupling between layer $l$ and $l^{\prime}$, where
$J_{\lambda,\lambda^{\prime}}^{l,l^{\prime}}$$\approx$ $\psi_{\lambda}^{l}\left(0\right)\left[\psi_{\lambda^{\prime}}^{l^{\prime}}\left(0\right)\right]^{\ast}\frac{V\left(\boldsymbol{k},\Delta z\right)}{4}\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda}^{+}\right)\times$
$\left(k_{+}e^{-i\theta_{l^{\prime}}}d_{cv,\lambda^{\prime}}^{-}+k_{-}e^{i\theta_{l^{\prime}}}d_{cv,\lambda^{\prime}}^{+}\right)^{\ast}.$
Here the Coulomb potential $V\left(\boldsymbol{k},\Delta z\right)=2\pi e^{2}/\left(\epsilon k\right)\times\exp\left(-k\Delta z\right)$ is related with the interlayer distance $\Delta z$ between two layers [39]. In the rest of this paper, we take $\Delta z=1\,\mathrm{nm}$ as a typical value. Since the optical transition dipole $d_{cv,\lambda}^{\gamma}$ can be got from the effective single-particle two-band $\boldsymbol{k}\cdot\boldsymbol{p}$ model near $\lambda$ valley, also the exciton Bohr radius $a_{B,\lambda}^{\gamma,l}$, the construction has been finished.
I.2 Excitonic Hamiltonians of TMDs
For the TMD cases, there are two inequivalent valleys as $\alpha\equiv\boldsymbol{K}$ and $\beta\equiv-\boldsymbol{K}$, connected by the time-reversal symmetry. It has been shown that the excitonic dispersion is isotropic for the first order of $\boldsymbol{k}$ in monolayer TMDs [13, 21, 23], these two properties lead to $H_{k}^{l}=\frac{\hbar^{2}k^{2}}{2m_{ex}}=H_{k}$ and $a_{B,\lambda}^{x,l}=a_{B,\lambda}^{y,l}=a_{B,\lambda^{\prime}}^{x,l^{\prime}}=a_{B,\lambda^{\prime}}^{y,l^{\prime}}=a_{B}$. Without loss of generality, we let $\theta_{b}=0$ and $\theta_{t}=\theta$ and replace the twist angle $\theta$ in momentum space into the real space one, i.e., $\theta\rightarrow-\theta$ in the rest of the paper. The effective Hamiltonian $H_{\mathrm{TMD}}$ of valley excitons in the twisted homobilayer TMD can be written as
$H_{\mathrm{TMD}}=\frac{\hbar^{2}k^{2}}{2m_{ex}}+\sum_{l=t,b}H_{\mathrm{intra}}^{l}+\sum_{l,l^{\prime}=t,b}H_{\mathrm{inter}}^{l,l^{\prime}},$ (17)
in the basis $\left\{\left|l,\boldsymbol{k}\right\rangle_{K},\left|l,\boldsymbol{k}\right\rangle_{-K}\right\}$, with
$H_{\mathrm{intra}}^{b}$ $=$ $J\frac{k}{K}\left(\begin{array}[]{cc}1&-e^{-2i\varphi}\\ -e^{2i\varphi}&1\end{array}\right),$
$H_{\mathrm{intra}}^{t}$ $=$ $J\frac{k}{K}\left(\begin{array}[]{cc}1&-e^{-2i\left(\theta+\varphi\right)}\\ -e^{2i\left(\theta+\varphi\right)}&1\end{array}\right),$
and
$H_{\mathrm{inter}}^{b,t}=\left(H_{\mathrm{inter}}^{t,b}\right)^{\dagger}=J\frac{k}{K}e^{-k\Delta z}\left(\begin{array}[]{cc}e^{i\theta}&-e^{-i\left(\theta+2\varphi\right)}\\ -e^{i\left(\theta+2\varphi\right)}&e^{-i\theta}\end{array}\right),$
giving the dispersion as [39]
$\varepsilon_{1}$ $=$ $\varepsilon_{2}=\frac{\hbar^{2}k^{2}}{2m_{ex}},$ (20)
$\varepsilon_{3}$ $=$ $\frac{\hbar^{2}k^{2}}{2m_{ex}}+2J\frac{k}{K}-2J\frac{k}{K}e^{-k\Delta z},$
$\varepsilon_{4}$ $=$ $\frac{\hbar^{2}k^{2}}{2m_{ex}}+2J\frac{k}{K}+2J\frac{k}{K}e^{-k\Delta z},$
which is independent of the twisted angle $\theta$. $m_{ex}\approx m_{e}$ is the effective exciton mass in monolayer TMDs. Here $K=4\pi/3a$, $a$ being TMD’s lattice constant, and $J\sim 1\,\mathrm{eV}$ can be extracted from first principle wavefunctions and exciton spectrum [7, 21, 23]. The isotropic properties of the excitonic bands in momentum space are evident from Fig. 1(c).
If there is a strain in the bottom layer of twisted bilayer TMDs, it can be transferred into the top layer with some loss [46, 47, 48]. This mechanism leads to chiral symmetric excitons in twisted bilayer TMDs, where the short-range e-h exchange contributes to the coupling between excitons in different valleys of the same layer. This results in the non-diagonal term of $H_{\mathrm{intra}}^{l}$ becoming $J_{0}^{l}-J\frac{k}{K}e^{\pm i\left(\theta_{l}+2\varphi\right)}$ [36]. In our consideration, $J_{0}^{b}\approx-6\,\mathrm{meV}$ arises from the strain along $x$ direction with an extent of $\epsilon_{st}=1\%$ [21], and $J_{0}^{t}\approx J_{0}^{b}\frac{\cos 2\theta+\sin 2\theta}{2}$ is twisted-angle dependent [36]. The related excitonic dispersions, with and without strain in untwisted situation, are displayed in Fig. 1(d), which also shows the polarization of the optical dipole for the band $\varepsilon_{n=1,3}$. The anisotropic polarization distribution of optical dipole in momentum space indicates that the chiral symmetry in real space preserves in the effective excitonic Hamiltonian when strain is applied.
I.3 Excitonic Hamiltonian of BP
The excitonic Hamiltonian of BP can be obtained in the basis of $\{\left|b,\boldsymbol{k}\right\rangle_{\Gamma},\left|t,\boldsymbol{k}\right\rangle_{\Gamma}\}$ by using the same approach as for TMDs. Since there is only one valley ($\Gamma$ point) per layer under the low-energy approximation, i.e., $\alpha=\beta\equiv\Gamma$, this Hamiltonian takes a simple $2\times 2$ form as [36]
$H_{\mathrm{BP}}$ $=$ $\left(\begin{array}[]{cc}\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex}^{\gamma,b}\right)&0\\ 0&\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex}^{\gamma,t}\right)\end{array}\right)$ (26)
$+\left(\begin{array}[]{cc}J_{\Gamma}^{b}&J_{\Gamma}^{b,t}\\ J_{\Gamma}^{t,b}&J_{\Gamma}^{t}\end{array}\right),$
The excitonic dispersions of $H_{\mathrm{BP}}$ has the from
$\varepsilon_{1,2}$ $=$ $\hbar^{2}\left(M_{x}^{+}k_{x}^{2}+M_{y}^{+}k_{y}^{2}\right)+$
$\frac{k}{X}\frac{\mathcal{J}^{b}\cos^{2}\varphi+\mathcal{J}^{t}\cos^{2}\left(\varphi+\theta\right)}{2}$
±$\pm\sqrt{A^{2}+B^{2}},$
where $A=\hbar^{2}\left(M_{x}^{-}k_{x}^{2}+M_{y}^{-}k_{y}^{2}\right)+\frac{k}{X}\frac{\mathcal{J}^{b}\cos^{2}\varphi-\mathcal{J}^{t}\cos^{2}\left(\varphi+\theta\right)}{2}$ and $B=\sqrt{\mathcal{J}^{b}\mathcal{J}^{t}}\frac{k}{X}e^{-kz}\cos\varphi\cos\left(\varphi+\theta\right)$ are functions of the twist angle $\theta$, with
$M_{x}^{\pm}=\frac{m_{ex}^{x,b}\pm m_{ex}^{x,t}}{4m_{ex}^{x,b}m_{ex}^{x,t}},M_{y}^{\pm}=\frac{m_{ex}^{y,b}\pm m_{ex}^{x,t}}{4m_{ex}^{y,b}m_{ex}^{y,t}}.$
Unlike in TMDs, $\theta$ affects the excitonic properties of $H_{\mathrm{BP}}$, which can be seen from the highly anisotropic dispersion at $\theta=0$ and $\pi/6$, also the highly anisotropic polarization distribution of optical dipole, as shown in Fig. 1(c).
Figure 2: (color online) (a) Schematic illustration of the time reversal even Hall effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ (Eq. 28) for $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) Similar plot for $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 2\,\mathrm{eV}$. $\tau=1\,\mathrm{ps}$ for all plots [43].
II The time-reversal even and crossed nonlinear dynamical Hall effect in those systems
II.1 The time-reversal even Hall effect (TREHE)
The non-zero TREHE requires chiral symmetry in real space [43], which is naturally satisfied by the twisted bilayer structure, as shown in Fig. 2(a). In twisted bilayer chiral excitonic systems, this effect can be defined as similar as the one for twisted bilayer electric systems [43], which comes from the semi-classical calculation for the current density of each layer
$\boldsymbol{j}^{\mathrm{sys/env}}=\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}\left(\boldsymbol{k}\right)\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right),$
where two layers are divided into the system layer (top layer) and environment layer (bottom layer). $\boldsymbol{j}^{\mathrm{sys}}=-\boldsymbol{j}^{\mathrm{env}}$ preserves by the Onsager relation [43]. $f_{n}\left(\boldsymbol{k}\right)\approx f_{n}^{0}\left(\boldsymbol{k}\right)-\tau\frac{df_{n}^{0}\left(\boldsymbol{k}\right)}{dt}$ is the off-equilibrium distribution function expanded in the first order of relaxation time $\tau$, with $f_{n}^{0}\left(\boldsymbol{k}\right)=1/\left[\exp\left(\frac{\varepsilon_{n}\left(\boldsymbol{k}\right)-\mu}{k_{B}T}\right)-1\right]$ is the equilibrium Bose-Einstein distribution function. $\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right)=\left\langle u_{n}\left(\boldsymbol{k}\right)\right|\frac{1}{2}\left\{\boldsymbol{\hat{v}},\hat{P}^{\mathrm{sys/env}}\right\}\left|u_{n}\left(\boldsymbol{k}\right)\right\rangle$ is the velocity projection on system/environment layer of band $n$, where $\hat{P}^{\mathrm{sys/env}}=\left(1\pm\hat{\sigma}_{z}\right)/2$ is the projection operator onto the system/environment layer, with the Pauli matrix $\hat{\sigma}_{z}$ being the out-of-plane layer-pseudospin (LPS) operator operating in the layer index subspace [49]. $\boldsymbol{\hat{v}}=\frac{\partial H_{\mathrm{ex}}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}$ is the total velocity operator.
If we consider that there is an inhomogeneous distribution of the temperature $T$ in real space and a general mechanical in-plane force $\boldsymbol{F}$ as $\boldsymbol{F}=\hbar\boldsymbol{\dot{k}}$. After some tedious derivations [36], we can get the time-reversal even Hall conductivity of the system as $\sigma_{H}^{\mathrm{sys}}=\tau\mathcal{V}/\hbar$, while $\alpha_{H}^{\mathrm{sys}}=\tau\mathcal{V}_{\mathrm{Ner}}/\hbar$ represents the time-reversal even Nernst conductivity, where
$\mathcal{V}$ $=$ $\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}^{0}\omega_{n}\left(\boldsymbol{k}\right),$ (28)
$\mathcal{V}_{\mathrm{Ner}}$ $=$ $\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}\frac{\varepsilon_{n}-\mu}{k_{B}T}f_{n}^{0}\omega_{n}\left(\boldsymbol{k}\right),$
with $\omega_{n}\left(\boldsymbol{k}\right)=\frac{1}{2}\left[\nabla_{\boldsymbol{k}}\times\boldsymbol{v}_{n}^{\mathrm{sys}}\left(\boldsymbol{k}\right)\right]$ means the $\boldsymbol{k}$-space vorticity of the layer current [43].
In TMDs without strain, although the chiral symmetry in real space is preserved, the rotational symmetry of the monolayer TMDs ensures that the effective excitonic Hamiltonian is independent of the twisted angle $\theta$ [39]. This reinstates the mirror symmetry and leads to isotropic dispersions, as shown in Fig. 1(d). Consequently, the TREHE is forbidden. However, if strain is applied to the bottom layer of twisted bilayer TMDs, strain transfer between the two layers [47, 48] will result in a twisted-angle-dependent strain in the top layer [47, 48, 36]. These two inequivalent strains in the different layers can maintain chiral symmetry in the excitonic Hamiltonian, leading to a obvious TREHE, as shown in Fig. 2(b). Both $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are periodic functions of $\theta$, with a period of $\pi/2$ due to $\theta$-dependent exchange coupling between excitons in different valleys of top layer, indicated by the product of two velocities in Fig. 3(a) [50].
In BP, due to the high anisotropy of the material, the chiral symmetry in real space is retained in the excitonic Hamiltonian, leading to anisotropic dispersions and a non-zero TREHE, as shown in Figs. 1(c) and 2(c). Here, $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are also periodic functions of $\theta$, but with a period of $\pi$ due to the $C_{2}$ symmetry of the product of two velocities within the integral, as illustrated in Fig. 3(b). The chemical potential $\mu$ in Fig. 2 is chosen such that $\min\varepsilon_{1}-\mu\approx 0.005\,\mathrm{eV}$, where $\varepsilon_{1}$ is the ground energy of the excitonic Hamiltonian, and $\min\varepsilon_{1}$ in both materials is independent of $\theta$. Based on this consideration, the changes of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ with respect to $T$ can be easily understand from the $T$ dependence of $f_{n=1}^{0}$ since it obviously increases as $T$ rises. This implies that the TREHE for excitons in chiral bilayer structures could be measurable at room temperature.
[…]
Figure 4: (color online) (a) Schematic illustration of the crossed nonlinear dynamical Hall effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}$ and $\alpha_{H}$ (Eq. 30) for effective excitonic Hamiltonian $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) The similar plot for effective excitonic Hamiltonian $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 1.98\,\mathrm{eV}$ with a gap $\varepsilon_{d}=0.02\,\mathrm{eV}$ induced by the physical control of layers such as strain. $F_{\perp}^{0}=0.1\,\mathrm{eV/nm}$, $\omega/2\pi=0.1\,\mathrm{THz}$ for all plots [44].
II.2 The crossed nonlinear dynamical Hall effect (CNDHE)
According to previous theoretical work [44, 45], this effect can be described from the similar starting point by calculating the current density of excitons in the intrinsic response
$\boldsymbol{j}=\sum_{n}\int\left[\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}^{0}\left(\boldsymbol{k}\right)\boldsymbol{v}_{n}\left(\boldsymbol{k}\right)+\nabla\times\boldsymbol{M}\left(r\right)\right],$ (29)
where
$\boldsymbol{v}_{n}\left(\boldsymbol{k}\right)=\frac{\partial\varepsilon_{n}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}+\dot{F}_{\perp}\boldsymbol{\Omega}_{n,F_{\perp}\boldsymbol{k}}-\frac{\boldsymbol{F}_{\parallel}}{\hbar}\times\Omega_{n,\boldsymbol{k}}\hat{z},$
is the velocity of an exciton in a bilayer system derived from semi-classical theory [44, 51, 52]. The velocity consists of three parts: the band velocity, the anomalous velocity induced by hybrid Berry curvature $\boldsymbol{\Omega}_{n,F_{\perp}\boldsymbol{k}}$ in the $\left(F_{\perp},\boldsymbol{k}\right)$ space, and by the $\boldsymbol{k}$-space Berry curvature $\Omega_{n,\boldsymbol{k}}$. Here $F_{\perp}=F_{\perp}^{0}f\left(t\right)$ is an out-of-plane mechanical force that may arises from the detuning of exciton resonance caused by layer dependent screening, with $f\left(t\right)$ being a function dependent on time $t$. $\boldsymbol{F}_{\parallel}=\hbar\frac{d\boldsymbol{k}}{dt}$ represents the in-plane general mechanical force, where
$\boldsymbol{M}\left(r\right)=\frac{k_{B}T}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}\Omega_{n,\boldsymbol{k}}\log\left(1-e^{-\left(\varepsilon_{n}-\mu\right)/k_{B}T}\right),$
is analogous to the equilibrium magnetization density in electronic systems [41, 53].
The semi-classical theory and symmetry analysis finally gives [36]
$\boldsymbol{j}$$\equiv$ $\boldsymbol{j}_{H}=\boldsymbol{j}^{\omega}\sin\left(\omega t\right),$ (30)
$\boldsymbol{j}^{\omega}$ $=$ $\boldsymbol{F}_{\parallel}\times\sigma_{H}\hat{z}+k_{B}\nabla T\times\alpha_{H}\hat{z},$
under an assumption that $F_{\perp}=F_{\perp}^{0}\cos\left(\omega t\right)$. The schematic illustration is shown in Fig. 4(a). $\sigma_{H}=\omega F_{\perp}^{0}\chi^{\mathrm{int}}$ is the crossed nonlinear dynamical Hall conductivity and $\alpha_{H}=\omega F_{\perp}^{0}\chi_{\mathrm{Ner}}^{\mathrm{int}}$ represents the crossed nonlinear dynamical Nernst conductivity, where
$\chi^{\mathrm{int}}$ $=$ $\frac{1}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{0}^{n}\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z},$
$\chi_{\mathrm{Ner}}^{\mathrm{int}}$ $=$ $\frac{1}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{0}^{n}\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z}\cdot$
$\left[\frac{\varepsilon_{n}-\mu}{k_{B}T}f_{n}^{0}-\log\left(f_{n}^{0}\exp\left(\frac{\varepsilon_{n}-\mu}{k_{B}T}\right)\right)\right],$
is intrinsic to the band structure, measuring the $\boldsymbol{k}$-space curl of the interlayer Berry connection polarizability (BCP)
$\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)=2\hbar^{2}\mathrm{Re}\sum_{m\neq n}\frac{\left\langle u_{n}\right|p_{z}\left|u_{m}\right\rangle}{\left(\varepsilon_{n}-\varepsilon_{m}\right)^{3}}\boldsymbol{v}_{mn},$
over the occupied states [44, 45]. Here, $p_{z}=\Delta z\hat{\sigma}_{z}$ is the interlayer dipole moment. Although CNDHE remains forbidden in TMDs without strain since the high symmetry of the related excitonic Hamiltonian, it can emerge after applying strain, as demonstrated in Fig. 4(b), which shows a similar dependence to that observed in the TREHE calculation. This behavior is dominated by the product of the band velocity and BCP, as shown in Fig. 3(c) [50]. In BP, however, a significant value of CNDHE appears at specific twisted angle $\theta$ (Fig. 4(c)), corresponding to the distribution of BCP with respect to $\theta$, as shown in Fig. 3(d). The periodicity dependent on $\theta$ is $\pi$ since only $C_{2}$ symmetry are left in $H_{\mathrm{BP}}$. Compared to the TREHE, the CNDHE is more dependent on the material’s anisotropy and the specific twisted angle $\theta$, as evidenced by the numerical calculations in Fig. 3(d). This finding may suggest new opportunities for applications of BP in valley exciton optoelectronics.
Acknowledgements.
C. L. would like to thank D. W. Zhai, B. Fu, and B. B for useful discussions. This work is supported by the National Key R&\&&D Program of China (2020YFA0309600), Research Grant Council of Hong Kong SAR (AoE/P-701/20, HKU SRFS2122-7S05), and New Cornerstone Science Foundation.