### v2

# Chiral excitonic systems in twisted bilayers from Förster coupling and unconventional excitonic Hall effects

Abstract
In twisted bilayer semiconductors with arbitrary twisting angles, a chiral excitonic system can arise from the interlayer electron-hole Coulomb exchange interaction (Förster coupling) that hybridizes the anisotropic intralayer excitons from individual layers. We present a general framework for the effective exciton Hamiltonian taking into account the electron-hole Coulomb exchange, using twisted homobilayer systems composed of transition metal dichalcogenides or black phosphorus as examples. We demonstrate that such chiral excitonic systems can feature unconventional Hall (Nernst) effects arising from quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry, for example, the time-reversal even layer Hall counter flow and the crossed nonlinear dynamical Hall effect, when mechanical and statistical force (temperature or density gradient) drives the exciton flow.
Layered two-dimensional (2D) semiconductors have provided a platform for the various frontiers of condensed matter physics [1, 2, 3, 4, 5, 6]. In particular, they offer exciting opportunities to explore exciton physics and strong light-matter interactions [2, 3, 4]. In transition metal dichalcogenides (TMDs), tightly bound Wannier excitons are formed around the degenerate ±$\pm K$ valleys at the corners of the Brillouin zone (BZ), where optical selection rules make possible manipulation of the valley pseudospin [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The small Bohr radius $\sim O(1)$ nm [17, 18, 19] underlies a significant electron-hole (e-h) Coulomb exchange at finite center-of-mass (COM) momentum [13, 20, 21, 22]. This effective coupling between the exciton’s valley pseudospin and COM degrees of freedom results in the splitting of the exciton dispersion into two branches having linearly polarized optical dipoles, longitudinal ($L$) and transverse ($T$) to exciton momentum, respectively [13, 21, 23, 24, 25, 26]. Recent studies have found exotic properties of excitons in monolayer TMDs on patterned substrates [27, 28] due to the massless $L$ branch’s sensitive dependence on the surrounding dielectric [29, 30]. Additionally, applying strain to monolayer TMDs can break the rotational symmetry. This symmetry breaking results in a non-zero contribution from short-range e-h exchange interactions between excitons in different valleys, leading to anisotropy and the formation of linearly dispersing Dirac saddle points within the light cone [13, 21].
Black phosphorus (BP) is an example of highly anisotropic 2D semiconductors. The monolayer has a direct bandgap of $1\sim 2\,\mathrm{eV}$ at the $\Gamma$ point in the BZ [31, 32, 33], suitable for hosting field-effect transistors as well [34]. The highly anisotropic band structure of the band edges leads to anisotropic intralayer excitons [32, 35], which exhibit linearly polarized optical dipoles, in contrast to the circularly polarized one for isotropic valley excitons in monolayer TMDs [21, 22, 23].
In layered structures, exciton’s layer degree of freedom can also be explored for novel functionalities, which is typically introduced by interlayer carrier tunneling that leads to the hybridization of intra- and inter-layer exciton species [3, 4, 21, 37]. On the other hand, Förster coupling [38], i.e., the e-h Coulomb exchange, provides another channel to introduce nontrivial layer structures for excitons. This interaction can non-locally transfer an intralayer exciton between different layers [39, 40], enabling exciton propagation in the out-of-plane direction. Its long-range nature allows it to function even when carrier interlayer tunneling is quenched by spacer layers. Previous studies have shown that cross-dimensional valley excitons can arise from this coupling in arbitrarily twisted stacks of monolayer semiconductors [41, 42].
Like electrons where the band geometric quantities have been extensively explored in transport phenomena [43], excitons can also have such properties from the dependence of its internal degree of freedom on the COM momentum, giving rise to novel exciton transport driven by the external mechanical and statistical force (temperature or density gradient) [44, 45]. In layered structures, this possibility can be rooted in the layer degree of freedom in the exciton wavefunction which can be engineered by twisting in the layered structure. Recently, two new types of Hall effect have been discovered in chiral electronic systems in twisted bilayers, i.e., the time-reversal even layer Hall counter flow (TREHCF) [46] and the crossed nonlinear dynamical Hall effect (CNDHE) [47, 48]. These findings uncover novel Hall physics in chiral electronic systems that can be straightforwardly engineered by twisting. The question that naturally arises is whether chiral excitonic systems can be similarly engineered.
In this paper, we show that the structural chirality from twisting in a layered structure can lead to a chiral exciton system when Förster coupling hybridizes intralayer excitons from the individual layers. In the limit where carrier interlayer tunneling is quenched, we present a general formulation of the effective exciton Hamiltonian taking into account the e-h Coulomb exchange interactions, using twisted homobilayer TMDs and BP as examples. The excitonic chirality requires the presence of anisotropy that breaks the continuous in-plane rotational symmetry, which is satisfied by the strong anisotropic exciton in BP. In the TMDs case, the required anisotropy can be provided by either trigonal warping or strain, and we considered the latter as an example. We demonstrate that such chiral excitonic systems can also feature unconventional Hall (Nernst) effects arising from quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry, for example, the TREHCF and the CNDHE, when mechanical and statistical force (temperature or density gradient) drives the longitudinal exciton flow. These findings point to exciton-based optoelectronic functionalities exploiting the layer quantum degree of freedom in twisted structures.
Figure 1: (color online) (a) Left panel: Schematic of an arbitrarily twisted homobilayer TMDs with strain (applied along $x$ direction with strain extent $\epsilon_{st}\sim 1\%$ in the bottom layer), where hBN spacers quench the charge hopping. Right panel: Illustration of electron-hole Coulomb exchange between valley excitons in the same or different layers, showing approximate band edges near high-symmetry points. (b) Similar plot for twisted homobilayer BP. (c) Left panel: Excitonic dispersions for the effective exciton Hamiltonian (17) of TMDs with and without strain in untwisted situation. Right panel: Polarization distribution of the optical dipole for two bands in momentum space. For clarity, only the polarization distribution $p_{n}$ of the optical dipole for the marked band $\varepsilon_{n}$ are shown. (d) Similar plots for the effective exciton Hamiltonian (26) of BP with different twisted angles $\theta$.
I Effective exciton Hamiltonians of twisted homobilayer TMDs and BP
As previously stated, the effective exciton Hamiltonian in twisted bilayer structures is primarily based on intralayer valley excitons in the two layers. In the case of monolayer situations under the low-energy approximation, there are typically at least two non-equivalent valleys connected by rotational or time-reversal symmetry in the first BZ, such as the ±$\pm K$ valley in monolayer TMDs [2, 7] (Fig. 1(a)), resulting in a $4\times 4$ effective Hamiltonian for intralayer excitons, as discussed below. However, when a single high symmetry point is present, such as the $\Gamma$ point in monolayer BP (Fig. 1(b)), the Hamiltonian reduces to a simple $2\times 2$ matrix. In the following, we present a method for obtaining the effective exciton Hamiltonian through a general calculation of the e-h Coulomb exchange interaction, and use it to acquire the exciton Hamiltonian for these two materials.
I.1 General description the effective exciton Hamiltonian
The construction of this Hamiltonian relies on two types of interaction: the e-h exchange interaction between intralayer excitons of the same or different valleys in the same layer and the Förster coupling that connects intralayer excitons in different layers, as illustrated in Fig. 1. For a specific layer $l$ ($l=t,b$ in this paper, $t/b$ means top/bottom layer, as shown in Fig. 1(a)), the basis of valley excitons can be expressed as $\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$, $\alpha,\beta$ means two different valley indexes with the in-plane COM momentum $\boldsymbol{k}=\left(k\cos\varphi,k\sin\varphi\right)$, and generally the effective exciton Hamiltonian in this situation can be expressed as
$H_{\mathrm{ex}}=\left(\begin{array}[]{cc}H_{k}^{b}&0\\ 0&H_{k}^{t}\end{array}\right)+\left(\begin{array}[]{cc}H_{\mathrm{intra}}^{b}&H_{\mathrm{inter}}^{b,t}\\ H_{\mathrm{inter}}^{t,b}&H_{\mathrm{intra}}^{t}\end{array}\right).$ (1)
The first term is the kinetic energy of excitons in different layers, which has the form
$H_{k}^{l}=\left(\begin{array}[]{cc}\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex,\alpha}^{\gamma,l}\right)&0\\ 0&\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex,\beta}^{\gamma,l}\right)\end{array}\right),$ (2)
with $m_{ex,\alpha}^{\gamma,l},\gamma=x,y$ is the exciton mass in $l$ layer, depending on the twisted angle $\theta_{l}$ of each layer (see Supplementary [36] for more details).
$\left(\begin{array}[]{c}\left\langle l,\boldsymbol{k}\right|_{\alpha}\\ \left\langle l,\boldsymbol{k}\right|_{\beta}\end{array}\right)H_{\mathrm{intra}}^{l}\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$ (5)
$\equiv$ $\left(\begin{array}[]{cc}J_{\alpha,\alpha}^{l}&J_{\alpha,\beta}^{l}\\ J_{\beta,\alpha}^{l}&J_{\beta,\beta}^{l}\end{array}\right),$ (8)
represents the e-h intra- or intervalley exchange between the excitons in the same or different valleys in the layer $l$, where
$J_{\lambda,\lambda^{\prime}}^{l}$$\approx$ $\left[\psi_{\lambda}^{l}\left(0\right)\right]^{\ast}\psi_{\lambda^{\prime}}^{l}\left(0\right)\frac{V\left(\boldsymbol{k}\right)}{4}\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda}^{+}\right)\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda^{\prime}}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda^{\prime}}^{+}\right)^{\ast},$
with $k_{\pm}=k_{x}\pm ik_{y}$, $d_{cv,\lambda}^{\pm}=d_{cv,\lambda}^{x}\pm id_{cv,\lambda}^{y}$. $d_{cv,\lambda/\lambda^{\prime}}^{\gamma},\lambda,\lambda^{\prime}=\alpha,\beta$ is the optical transition dipole in $\gamma$ direction, between conduction ($c$) and valence ($v$) band edges of $\lambda$ or $\lambda^{\prime}$ valley. $\psi_{\lambda}^{l}\left(0\right)\sim 1/\sqrt{a_{B,\lambda}^{x,l}a_{B,\lambda}^{y,l}}$ with the exciton Bohr radius $a_{B,\lambda}^{\gamma,l}$ for $\lambda$ valley can be seen as the square root of the probability for electron and hole to overlap in an exciton. $V\left(\boldsymbol{k}\right)=2\pi e^{2}/\left(\epsilon k\right)$ is the unscreened form for the Coulomb potential with $\epsilon\equiv 4\pi\epsilon_{0}\epsilon_{r}$ [13, 21, 23].
$\left(\begin{array}[]{c}\left\langle l,\boldsymbol{k}\right|_{\alpha}\\ \left\langle l,\boldsymbol{k}\right|_{\beta}\end{array}\right)H_{\mathrm{inter}}^{l,l^{\prime}}\left(\left|l^{\prime},\boldsymbol{k}\right\rangle_{\alpha},\left|l^{\prime},\boldsymbol{k}\right\rangle_{\beta}\right)$ (12)
$\equiv$ $\left(\begin{array}[]{cc}J_{\alpha,\alpha}^{l,l^{\prime}}&J_{\alpha,\beta}^{l,l^{\prime}}\\ J_{\beta,\alpha}^{l,l^{\prime}}&J_{\beta,\beta}^{l,l^{\prime}}\end{array}\right),$ (15)
shows the Förster coupling between layer $l$ and $l^{\prime}$, where
$J_{\lambda,\lambda^{\prime}}^{l,l^{\prime}}$$\approx$ $\psi_{\lambda}^{l}\left(0\right)\left[\psi_{\lambda^{\prime}}^{l^{\prime}}\left(0\right)\right]^{\ast}\frac{V\left(\boldsymbol{k},\Delta z\right)}{4}\times$
$\left(k_{+}e^{-i\theta_{l}}d_{cv,\lambda}^{-}+k_{-}e^{i\theta_{l}}d_{cv,\lambda}^{+}\right)\times$
$\left(k_{+}e^{-i\theta_{l^{\prime}}}d_{cv,\lambda^{\prime}}^{-}+k_{-}e^{i\theta_{l^{\prime}}}d_{cv,\lambda^{\prime}}^{+}\right)^{\ast}.$
Here the Coulomb potential $V\left(\boldsymbol{k},\Delta z\right)=2\pi e^{2}/\left(\epsilon k\right)\times\exp\left(-k\Delta z\right)$ is related with the interlayer distance $\Delta z$ between two layers [41]. In the rest of this paper, we take $\Delta z=1\,\mathrm{nm}$ as a typical value. Since the optical transition dipole $d_{cv,\lambda}^{\gamma}$ can be got from the effective single-particle two-band $\boldsymbol{k}\cdot\boldsymbol{p}$ model near $\lambda$ valley, also the exciton Bohr radius $a_{B,\lambda}^{\gamma,l}$, the construction has been finished.
I.2 The exciton Hamiltonian of TMDs
For the TMD cases, there are two inequivalent valleys as $\alpha\equiv\boldsymbol{K}$ and $\beta\equiv-\boldsymbol{K}$, connected by the time-reversal symmetry. It has been shown that the exciton dispersion is isotropic for the first order of $\boldsymbol{k}$ in monolayer TMDs [13, 21, 23], these two properties lead to $H_{k}^{l}=\frac{\hbar^{2}k^{2}}{2m_{ex}}=H_{k}$ and $a_{B,\lambda}^{x,l}=a_{B,\lambda}^{y,l}=a_{B,\lambda^{\prime}}^{x,l^{\prime}}=a_{B,\lambda^{\prime}}^{y,l^{\prime}}=a_{B}$. Without loss of generality, we let $\theta_{b}=0$ and $\theta_{t}=\theta$ and replace the twist angle $\theta$ in momentum space into the real space one, i.e., $\theta\rightarrow-\theta$ in the rest of the paper. The effective Hamiltonian $H_{\mathrm{TMD}}$ of valley excitons in the twisted homobilayer TMD can be written as
$H_{\mathrm{TMD}}=\frac{\hbar^{2}k^{2}}{2m_{ex}}+\sum_{l=t,b}H_{\mathrm{intra}}^{l}+\sum_{l,l^{\prime}=t,b}H_{\mathrm{inter}}^{l,l^{\prime}},$ (17)
in the basis $\left\{\left|l,\boldsymbol{k}\right\rangle_{K},\left|l,\boldsymbol{k}\right\rangle_{-K}\right\}$, with
$H_{\mathrm{intra}}^{b}$ $=$ $J\frac{k}{K}\left(\begin{array}[]{cc}1&-e^{-2i\varphi}\\ -e^{2i\varphi}&1\end{array}\right),$
$H_{\mathrm{intra}}^{t}$ $=$ $J\frac{k}{K}\left(\begin{array}[]{cc}1&-e^{-2i\left(\theta+\varphi\right)}\\ -e^{2i\left(\theta+\varphi\right)}&1\end{array}\right),$
and
$H_{\mathrm{inter}}^{b,t}=\left(H_{\mathrm{inter}}^{t,b}\right)^{\dagger}=J\frac{k}{K}e^{-k\Delta z}\left(\begin{array}[]{cc}e^{i\theta}&-e^{-i\left(\theta+2\varphi\right)}\\ -e^{i\left(\theta+2\varphi\right)}&e^{-i\theta}\end{array}\right),$
giving the dispersion as [41]
$\varepsilon_{1}$ $=$ $\varepsilon_{2}=\frac{\hbar^{2}k^{2}}{2m_{ex}},$ (20)
$\varepsilon_{3}$ $=$ $\frac{\hbar^{2}k^{2}}{2m_{ex}}+2J\frac{k}{K}-2J\frac{k}{K}e^{-k\Delta z},$
$\varepsilon_{4}$ $=$ $\frac{\hbar^{2}k^{2}}{2m_{ex}}+2J\frac{k}{K}+2J\frac{k}{K}e^{-k\Delta z},$
which is independent of the twisted angle $\theta$. $m_{ex}\approx m_{e}$ is the effective exciton mass in monolayer TMDs. Here $K=4\pi/3a$, $a$ being TMD’s lattice constant, and $J\sim 1\,\mathrm{eV}$ can be extracted from first principle wavefunctions and exciton spectrum [7, 21, 23]. The isotropic properties of the exciton bands in momentum space are evident from Fig. 1(c).
If there is a strain in the bottom layer of twisted bilayer TMDs, it can be transferred into the top layer with some loss [49, 50, 51]. This mechanism leads to chiral symmetric excitons in twisted bilayer TMDs, where the short-range e-h exchange contributes to the coupling between excitons in different valleys of the same layer. This results in the non-diagonal term of $H_{\mathrm{intra}}^{l}$ becoming $J_{0}^{l}-J\frac{k}{K}e^{\pm i\left(\theta_{l}+2\varphi\right)}$ [36]. In our consideration, $J_{0}^{b}\approx-6\,\mathrm{meV}$ arises from the strain along $x$ direction with an extent of $\epsilon_{st}=1\%$ [21], and $J_{0}^{t}\approx J_{0}^{b}\frac{\cos 2\theta+\sin 2\theta}{2}$ is twisted-angle dependent [36]. The related exciton dispersions, with and without strain in untwisted situation, are displayed in Fig. 1(c), which also shows the polarization of the optical dipole for the band $\varepsilon_{n=1,3}$. The anisotropic polarization distribution of optical dipole in momentum space indicates that the excitonic chirality appears in twisted bilayer TMDs when strain is applied.
I.3 The exciton Hamiltonian of BP
The exciton Hamiltonian of BP can be obtained in the basis of $\{\left|b,\boldsymbol{k}\right\rangle_{\Gamma},\left|t,\boldsymbol{k}\right\rangle_{\Gamma}\}$ by using the same approach as for TMDs. Since the band edge only appear in the $\Gamma$ point per layer under the low-energy approximation, i.e., $\alpha=\beta\equiv\Gamma$, this Hamiltonian takes a simple $2\times 2$ form as [36]
$H_{\mathrm{BP}}$ $=$ $\left(\begin{array}[]{cc}\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex}^{\gamma,b}\right)&0\\ 0&\sum_{\gamma}\hbar^{2}k_{\gamma}^{2}/\left(2m_{ex}^{\gamma,t}\right)\end{array}\right)$ (26)
$+\left(\begin{array}[]{cc}J_{\Gamma}^{b}&J_{\Gamma}^{b,t}\\ J_{\Gamma}^{t,b}&J_{\Gamma}^{t}\end{array}\right),$
The exciton dispersions of $H_{\mathrm{BP}}$ has the from
$\varepsilon_{1,2}$ $=$ $\hbar^{2}\left(M_{x}^{+}k_{x}^{2}+M_{y}^{+}k_{y}^{2}\right)+$
$\frac{k}{X}\frac{\mathcal{J}^{b}\cos^{2}\varphi+\mathcal{J}^{t}\cos^{2}\left(\varphi+\theta\right)}{2}$
±$\pm\sqrt{A^{2}+B^{2}},$
where $A=\hbar^{2}\left(M_{x}^{-}k_{x}^{2}+M_{y}^{-}k_{y}^{2}\right)+\frac{k}{X}\frac{\mathcal{J}^{b}\cos^{2}\varphi-\mathcal{J}^{t}\cos^{2}\left(\varphi+\theta\right)}{2}$ and $B=\sqrt{\mathcal{J}^{b}\mathcal{J}^{t}}\frac{k}{X}e^{-kz}\cos\varphi\cos\left(\varphi+\theta\right)$ are functions of the twist angle $\theta$, with
$M_{x}^{\pm}=\frac{m_{ex}^{x,b}\pm m_{ex}^{x,t}}{4m_{ex}^{x,b}m_{ex}^{x,t}},M_{y}^{\pm}=\frac{m_{ex}^{y,b}\pm m_{ex}^{x,t}}{4m_{ex}^{y,b}m_{ex}^{y,t}}.$
Unlike in TMDs, $\theta$ affects the excitonic properties of $H_{\mathrm{BP}}$, which can be seen from the highly anisotropic dispersion at $\theta=0$ and $\pi/6$, also the highly anisotropic polarization distribution of optical dipole, as shown in Fig. 1(d).
Figure 2: (color online) (a) Schematic illustration of the time reversal even layer Hall (Nernst) counter flow for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ (Eq. 28) for $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) Similar plot for $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 2\,\mathrm{eV}$. $\tau=1\,\mathrm{ps}$ for all plots [46].
II The time-reversal even layer Hall (Nernst) counter flow and crossed nonlinear dynamical Hall (Nernst) effect
II.1 The time-reversal even layer Hall (Nernst) counter flow (TREHCF)
The non-zero TREHCF requires chiral symmetry in real space [46], which is naturally satisfied by the twisted bilayer structure, as shown in Fig. 2(a). In twisted bilayer chiral excitonic systems, this effect can be defined as similar as the one for twisted bilayer electric systems [46], which comes from the semi-classical calculation for the current density of each layer
$\boldsymbol{j}^{\mathrm{sys/env}}=\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}\left(\boldsymbol{k}\right)\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right),$
where two layers are divided into the system layer (top layer) and environment layer (bottom layer). $\boldsymbol{j}^{\mathrm{sys}}=-\boldsymbol{j}^{\mathrm{env}}$ preserves by the Onsager relation [46]. $f_{n}\left(\boldsymbol{k}\right)\approx f_{n}^{0}\left(\boldsymbol{k}\right)-\tau\frac{df_{n}^{0}\left(\boldsymbol{k}\right)}{dt}$ is the off-equilibrium distribution function expanded in the first order of relaxation time $\tau$, with $f_{n}^{0}\left(\boldsymbol{k}\right)=1/\left[\exp\left(\frac{\varepsilon_{n}\left(\boldsymbol{k}\right)-\mu}{k_{B}T}\right)-1\right]$ is the equilibrium Bose-Einstein distribution function. $\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right)=\left\langle u_{n}\left(\boldsymbol{k}\right)\right|\frac{1}{2}\left\{\boldsymbol{\hat{v}},\hat{P}^{\mathrm{sys/env}}\right\}\left|u_{n}\left(\boldsymbol{k}\right)\right\rangle$ is the velocity projection on system/environment layer of band $n$, where $\hat{P}^{\mathrm{sys/env}}=\left(1\pm\hat{\sigma}_{z}\right)/2$ is the projection operator onto the system/environment layer, with the Pauli matrix $\hat{\sigma}_{z}$ being the out-of-plane layer-pseudospin operator operating in the layer index subspace [52]. $\boldsymbol{\hat{v}}=\frac{\partial H_{\mathrm{ex}}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}$ is the total velocity operator.
If we consider that there is an inhomogeneous distribution of the temperature $T$ in real space and a general mechanical in-plane force $\boldsymbol{F}$ as $\boldsymbol{F}=\hbar\boldsymbol{\dot{k}}$. After some tedious derivations [36], we can get the time-reversal even Hall conductivity for the layer counter flow of the system as $\sigma_{H}^{\mathrm{sys}}=\tau\mathcal{V}/\hbar$, while $\alpha_{H}^{\mathrm{sys}}=\tau\mathcal{V}_{\mathrm{Ner}}/\hbar$ represents the time-reversal even Nernst conductivity for the layer counter flow, where
$\mathcal{V}$ $=$ $\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}^{0}\omega_{n}\left(\boldsymbol{k}\right),$ (28)
$\mathcal{V}_{\mathrm{Ner}}$ $=$ $\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}\frac{\varepsilon_{n}-\mu}{k_{B}T}f_{n}^{0}\omega_{n}\left(\boldsymbol{k}\right),$
with $\omega_{n}\left(\boldsymbol{k}\right)=\frac{1}{2}\left[\nabla_{\boldsymbol{k}}\times\boldsymbol{v}_{n}^{\mathrm{sys}}\left(\boldsymbol{k}\right)\right]$ means the $\boldsymbol{k}$-space vorticity of the layer current [46], exhibiting the nontrivial quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry.
In TMDs without strain, although the chiral symmetry in real space is preserved, the rotational symmetry of the monolayer TMDs ensures that the effective exciton Hamiltonian is independent of the twisted angle $\theta$ [41]. This reinstates the mirror symmetry and leads to isotropic dispersions, as shown in Fig. 1(c). Consequently, the TREHCF is forbidden. However, if strain is applied to the bottom layer of twisted bilayer TMDs, strain transfer between the two layers [50, 51] will result in a twisted-angle-dependent strain in the top layer [50, 51, 36]. These two inequivalent strains in the different layers can break the rotational symmetry in the exciton Hamiltonian, leading to the excitonic chirality and a obvious TREHCF, as shown in Fig. 2(b). Both $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are periodic functions of $\theta$, with a period of $\pi/2$ dominated by the quantum geometric properties of $\omega_{n}\left(\boldsymbol{k}\right)$, as shown from the product of two velocities in Fig. 3(a) [53].
In BP, due to the high anisotropy, the excitonic chirality is naturally satisfied, leading to anisotropic exciton dispersions and a non-zero TREHCF, as shown in Figs. 1(d) and 2(c). Here, $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are also periodic functions of $\theta$, but with a period of $\pi$ due to the $C_{2}$ symmetry of the product of two velocities within the integral, as illustrated in Fig. 3(b). The chemical potential $\mu$ in Fig. 2 is chosen such that $\min\varepsilon_{1}-\mu\approx 0.005\,\mathrm{eV}$, where $\varepsilon_{1}$ is the ground energy of the exciton Hamiltonian, and $\min\varepsilon_{1}$ in both materials is independent of $\theta$. Based on this consideration, the changes of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ with respect to $T$ can be easily understand from the $T$ dependence of $f_{n=1}^{0}$ since it obviously increases as $T$ rises. This implies that the TREHCF for excitons in chiral bilayer structures could be measurable at room temperature.
[…]
Figure 4: (color online) (a) Schematic illustration of the crossed nonlinear dynamical Hall (Nernst) effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}$ and $\alpha_{H}$ (Eq. 30) for effective exciton Hamiltonian $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) The similar plot for effective exciton Hamiltonian $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 1.98\,\mathrm{eV}$ with a gap $\varepsilon_{d}=0.02\,\mathrm{eV}$ induced by the physical control of layers such as strain. $F_{\perp}^{0}=0.1\,\mathrm{eV/nm}$, $\omega/2\pi=0.1\,\mathrm{THz}$ for all plots [47].
II.2 The crossed nonlinear dynamical Hall (Nernst) effect (CNDHE)
According to previous theoretical works [47, 48], this effect can be described from the similar starting point by calculating the current density of excitons in the intrinsic response
$\boldsymbol{j}=\sum_{n}\int\left[\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{n}^{0}\left(\boldsymbol{k}\right)\boldsymbol{v}_{n}\left(\boldsymbol{k}\right)+\nabla\times\boldsymbol{M}\left(r\right)\right],$ (29)
where
$\boldsymbol{v}_{n}\left(\boldsymbol{k}\right)=\frac{\partial\varepsilon_{n}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}+\dot{F}_{\perp}\boldsymbol{\Omega}_{n,F_{\perp}\boldsymbol{k}}-\frac{\boldsymbol{F}_{\parallel}}{\hbar}\times\Omega_{n,\boldsymbol{k}}\hat{z},$
is the velocity of an exciton in a bilayer system derived from semi-classical theory [47, 54, 55]. The velocity consists of three parts: the band velocity, the anomalous velocity induced by hybrid Berry curvature $\boldsymbol{\Omega}_{n,F_{\perp}\boldsymbol{k}}$ in the $\left(F_{\perp},\boldsymbol{k}\right)$ space, and by the $\boldsymbol{k}$-space Berry curvature $\Omega_{n,\boldsymbol{k}}$. Here $F_{\perp}=F_{\perp}^{0}f\left(t\right)$ is an out-of-plane mechanical force that may arises from the detuning of exciton resonance caused by layer dependent screening, with $f\left(t\right)$ being a function dependent on time $t$. $\boldsymbol{F}_{\parallel}=\hbar\frac{d\boldsymbol{k}}{dt}$ represents the in-plane general mechanical force, where
$\boldsymbol{M}\left(r\right)=\frac{k_{B}T}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}\Omega_{n,\boldsymbol{k}}\log\left(1-e^{-\left(\varepsilon_{n}-\mu\right)/k_{B}T}\right),$
is analogous to the equilibrium magnetization density in electronic systems [44, 56].
The semi-classical theory and symmetry analysis finally gives [36]
$\boldsymbol{j}$$\equiv$ $\boldsymbol{j}_{H}=\boldsymbol{j}^{\omega}\sin\left(\omega t\right),$ (30)
$\boldsymbol{j}^{\omega}$ $=$ $\boldsymbol{F}_{\parallel}\times\sigma_{H}\hat{z}+k_{B}\nabla T\times\alpha_{H}\hat{z},$
under an assumption that $F_{\perp}=F_{\perp}^{0}\cos\left(\omega t\right)$. The schematic illustration is shown in Fig. 4(a). $\sigma_{H}=\omega F_{\perp}^{0}\chi^{\mathrm{int}}$ is the crossed nonlinear dynamical Hall conductivity and $\alpha_{H}=\omega F_{\perp}^{0}\chi_{\mathrm{Ner}}^{\mathrm{int}}$ represents the crossed nonlinear dynamical Nernst conductivity, where
$\chi^{\mathrm{int}}$ $=$ $\frac{1}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{0}^{n}\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z},$
$\chi_{\mathrm{Ner}}^{\mathrm{int}}$ $=$ $\frac{1}{\hbar}\sum_{n}\int\frac{d^{2}\boldsymbol{k}}{\left(2\pi\right)^{2}}f_{0}^{n}\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z}\cdot$
$\left[\frac{\varepsilon_{n}-\mu}{k_{B}T}f_{n}^{0}-\log\left(f_{n}^{0}\exp\left(\frac{\varepsilon_{n}-\mu}{k_{B}T}\right)\right)\right],$
is intrinsic to the band structure. Here, the unusual quantum geometric properties induced by the chiral symmetric layer hybridized wavefunctions is measured by the $\boldsymbol{k}$-space curl of the interlayer Berry connection polarizability (BCP)
$\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)=2\hbar^{2}\mathrm{Re}\sum_{m\neq n}\frac{\left\langle u_{n}\right|p_{z}\left|u_{m}\right\rangle}{\left(\varepsilon_{n}-\varepsilon_{m}\right)^{3}}\boldsymbol{v}_{mn},$
over the occupied states [47, 48]. Here, $p_{z}=\Delta z\hat{\sigma}_{z}$ is the interlayer dipole moment. Although CNDHE remains forbidden in TMDs without strain since the rotational symmetry of the related exciton Hamiltonian, it can emerge after applying strain, as demonstrated in Fig. 4(b), which shows a similar dependence to that observed in the TREHCF calculation. This behavior is dominated by the quantum geometric properties of $\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z}$, as shown in Fig. 3(c) [53]. In BP, however, a significant value of CNDHE appears at specific twisted angle $\theta$ (Fig. 4(c)), corresponding to the distribution of BCP with respect to $\theta$, as shown in Fig. 3(d). The periodicity dependent on $\theta$ is $\pi$ since only $C_{2}$ symmetry are left in $H_{\mathrm{BP}}$. Compared to the TREHCF, the CNDHE is more dependent on the material’s anisotropy and the specific twisted angle $\theta$, as evidenced by the numerical calculations in Fig. 3(d). This finding may suggest new opportunities for applications of BP in exciton-based optoelectronics.
Acknowledgements.
C. L. would like to thank D. W. Zhai, B. Fu, C. Xiao, and B. B for useful discussions. This work is supported by the National Key R&\&&D Program of China (2020YFA0309600), Research Grant Council of Hong Kong SAR (AoE/P-701/20, HKU SRFS2122-7S05), and New Cornerstone Science Foundation.